# Feature Importance-Aware Deep Joint Source-Channel Coding for Computationally Efficient and Adjustable Image Transmission

### v3

@@ III-D Training and Loss Function Design
$z^{l}_{{\textrm{soft}}}(i,j)$ $=\frac{\exp{\left(\log\pi^{l}(i,j)+g^{l}(i,j)\right)/\tau}}{\sum_{m=1}^{2}\exp{\left(\log\pi^{m}(i,j)+g^{m}(i,j)\right)/\tau}},$
$\mathbf{z}_{{\textrm{hard}}}(i,j)$ $={\textrm{one\_hot}}\left(\operatorname*{arg\,max}_{l}\left[\log\pi^{l}(i,j)+g^{l}(i,j)\right]\right).$
@@
$\mathbf{z}_{{\textrm{ST}}}=\mathbf{z}^{1}_{{\textrm{hard}}}+\mathbf{z}^{1}_{{\textrm{soft}}}-{\textrm{stop\_grad}}(\mathbf{z}^{1}_{{\textrm{soft}}}).$
@@
$\mathbf{z}_{{\textrm{out}}}={\textrm{self\_attention}}({\textrm{deformation}}(\mathbf{z}_{{\textrm{ST}}}\odot_{\textrm{w}}\mathbf{z}_{{\textrm{in}}})).$
@@
$\bar{\mathbf{z}}^{e}_{{\textrm{ST}}}$ $=\frac{1}{N}\sum_{i=1}^{N}\bar{\mathbf{z}}^{e,i}_{{\textrm{ST}}},~~\bar{\mathbf{z}}^{d}_{{\textrm{ST}}}=\frac{1}{N}\sum_{i=1}^{N}\bar{\mathbf{z}}^{d,i}_{{\textrm{ST}}}~~\gamma_{{\textrm{train}}}<1,$
@@ IV-A Experimental Setting
To differentiate between these baselines, we refer to the initial deepJSCC [5] based on the convolution blocks and the deepJSCC based on residual convolution blocks [6] as ConvJSCC and ResJSCC, respectively. In contrast to SwinJSCC [7] that applies self-attention to all features, LICRFJSCC [29] applies self-attention only to a fixed subset of feature channels. Unless explicitly specified otherwise, the GFLOPs and memory usages of implemented models follow Table I. The scaling settings for ConvJSCC and ResJSCC follow those in [5], while the scaling configuration for SwinJSCC is determined via an exhaustive search to achieve the best possible performance under comparable GFLOPs constraints. LICRFJSCC, LAJSCC, and FAJSCC scale settings are based on SwinJSCC. The detailed implementation code is available at github.com/hansung-choi/FAJSCCv2.
@@
Datasets and Training To train models, we use the DIV2K training dataset [49], which consists of $800$ high-resolution images. To maintain consistent image resolution within each batch, we randomly crop images to a resolution of $128\times 128$ with a batch size of $32$ during training. The image distortion loss $\mathcal{L}_{{\textrm{distortion}}}$ is set to MSE and $1-{\textrm{SSIM}}$ when evaluating models for PSNR and SSIM, respectively. Our FAJSCC is trained to minimize $\mathcal{L}_{{\textrm{FA}}}$ in (1) while others are trained to minimize $\mathcal{L}_{{\textrm{distortion}}}$. The hyperparameter settings for $\mathcal{L}_{{\textrm{FA}}}$ are $\eta=1.0$, $\gamma_{{\textrm{train}}}=0.5$ , and $\tau=1$. Note that $\gamma_{{\textrm{train}}}$ is the target average importance ratio in the training phase; the importance ratio during the test phase could be different, e.g., Figure 9. All models are trained using Adam optimizer with a learning rate $0.0001$ for $200$ epochs.
We consider four CPP values $\frac{1}{12},\frac{1}{16},\frac{1}{24},\frac{1}{32}]$ and four SNR values [1,4,7,10][1,4,7,10]dB under AWGN and fast Rayleigh fading channels. For all models, the training SNR is the same as the test SNR. Since we observed unstable updates in ConvJSCC and ResJSCC during training with small CPP values, we selected the best-performing model out of five trained models for each to ensure reliable results. Test results are obtained from the DIV2K validation dataset [49] for numerical evaluation and the Kodak dataset [50] for visual inspection in Figure 10.
@@
Figure 6: Efficiency comparison for various model sizes with respect to computational burden (GFLOPs) and memory usage.
Figure 7: PSNR results under different channel and CPP environments.
Figure 8: SSIM results under different channel and CPP environments.
Figure 9: Performance comparison for various importance ratios under different channel noises. Varying $\gamma_{e}$: Only the importance ratio for the encoder’s FA blocks is changed. $\gamma_{e}=\gamma_{d}$: The importance ratios for the encoder’s and decoder’s FA blocks are changed together with the same value. Varying $\gamma_{d}$: Only the importance ratio for the decoder’s FA blocks is changed.
@@ IV-B Main Result
TABLE I: Comparison for GFLOPs and memory usage of deepJSCCs under AWGN channel, SNR=10dB10\ $\mathrm${dB}, and CPP=$1/32$.
PSNR $\mathrm{(dB)}$ GFLOPs Memory (MB)
ConvJSCC [5] 21.09 4.48 0.64
ResJSCC [6] 24.05 5.73 0.77
SwinJSCC [7] 25.90 20.02 71.26
LICRFJSCC [29] 24.52 9.02 3.84
LAJSCC (ours) 26.55 11.35 5.27
FAJSCC (ours) 26.86 17.58 13.56
Figure 10: The first row shows the transmitted image at ${\textrm{CPP}}=\frac{1}{12}$, ${\textrm{SNR}}=10$dB under the AWGN channel, and the third row shows the transmitted image at ${\textrm{CPP}}=\frac{1}{12}$, ${\textrm{SNR}}=1$dB under the AWGN channel. The second and fourth rows show the detailed results of the green box in the original images of the first and third rows.
Architecture Efficiency: Table I shows PSNR performance in terms of computational cost (GFLOPs) and memory usage (MB) under AWGN channel, SNR=1010dB, and CPP=$1/32$. Here, GFLOPs are measured when transmitting $512\times 768$ resolution image. In this table, our LAJSCC and FAJSCC achieve the highest PSNR while using smaller computational resources and memory than SwinJSCC [7].
To fully verify the performance gain of our model, we provide efficiency comparison for various model sizes in Figure 6. To match the computational resource usage for comparison, we implement smaller and larger versions of SwinJSCC, LAJSCC, and FAJSCC by decreasing and increasing feature sizes by about $50\%$. In the case of ConvJSCC and ResJSCC, we increase the feature size by factors of $2$ and $4$ to construct larger models.
Figure 6 shows the tradeoff between computational block complexity (GFLOPs and memory usage) and PSNR performance for various JSCC schemes. Most existing works, such as ConvJSCC [5], ResJSCC [6], and SwinJSCC [7], primarily aim to improve PSNR. These performance gains are achieved by employing heavier computational blocks, which consequently incur substantially higher GFLOPs and memory requirements. In particular, SwinJSCC demands approximately an order of magnitude more memory than ConvJSCC and ResJSCC, and imposes a significantly higher minimum computational burden for the same feature scales due to its reliance on heavy multi-head self-attention mechanisms. This increased requirement gives a large burden to practical deployment, especially for IoT devices. To address this issue, recent works decrease computational burden at the expense of performance degradation. For example, LICRFJSCC [29] applies advanced self-attention only to a fixed subset of feature channels without considering feature importance. Although this method reduces the minimum level of computational burden, it also compromises performance by shrinking the feature representation space.
@@
PSNR and SSIM Results: Figures 7 and 8 show the PSNR and SSIM performance results across various communication environments. Our LAJSCC and FAJSCC outperform the previous models for different performance metrics under various communication environments. This performance gain primarily comes from our dimension-specialized computation approaches: (1) Using only axis-dimension specialized operations (LAJSCC) shows higher performance than the recent SOTA SwinJSCC in most communication environments, while requiring computational resources about half and $14$ times smaller memory. This is a notable achievement, as even the recent SOTA lightweight model LICRFJSCC [29] reduces computational resources by shrinking feature sizes when applying self-attention, which inherently limits its performance, making it impossible to reach that of SwinJSCC. (2) Upon the axis-dimension specialized operations, our proposed selective deformable self-attention (FAJSCC) even more enhances the image transmission performance by utilizing deformation and feature importance. These advantages lead FAJSCC to outperform SwinJSCC while using smaller computational and memory resources.
@@
Computation Resource Adjustment: Our FAJSCC is the first deepJSCC model capable of independently adjusting computational resources of the encoder and decoder. It is important to note that typical previous deepJSCC models have fixed and symmetric computational complexity at encoder and decoder [5, 7, 12] or the decoder’s computational complexity is inherently constrained by feature deletion at the encoder [30, 34]. As a result, such independent computational resource adjustment has not been possible in previous models. On the contrary, FAJSCC achieves this adjustment by independently varying the importance ratio $\gamma_{e}$ for the encoder’s FA blocks and $\gamma_{d}$ for the decoder’s FA blocks, without needing to agree with each other’s value. Since available computational resources for the encoder and decoder may vary over time, this flexible adjustment mechanism enhances the practicality of FAJSCC for real-world communication applications.
@@
Complexity Demands at Encoder and Decoder: Recalling that deepJSCC involves the joint optimization of source coding and channel coding, we can further investigate the image reconstruction performance from three different perspectives. The first perspective is the encoder for source and channel coding where the extracted features should be both concise and robust to channel noise. The second perspective is the decoder for source coding where images are regenerated. The final perspective is the decoder for channel coding, which involves perceiving the meaning of the noisy received signal. By considering these components, our experiments with varying values of $\gamma_{e}$ and $\gamma_{d}$ reveal the relationship between computational complexity and the contribution of each component to reconstruction performance. It is important to note that previous deepJSCC models have fixed computational complexity at the encoder and decoder [5, 7, 12] or the decoder’s computational complexity is inherently constrained by feature deletion at the encoder [30, 34]. As a result, such a discussion has not been possible in previous models.
First, when varying $\gamma_{e}$ (triangles in Figure 9), no significant change in PSNR is observed at both ${\textrm{SNR}}=1$dB (low) and ${\textrm{SNR}}=10$dB (high). This indicates that robust and concise feature extraction in the encoder does not require substantial computational resources, and adjusting the encoder’s computational budget has minimal impact on overall communication performance. On the other hand, when varying $\gamma_{d}$ (crosses in Figure 9), PSNR exhibits different behaviors depending on the SNR. The PSNR variation at ${\textrm{SNR}}=1$dB is approximately 0.280.28dB, which is about three times of the 0.090.09dB variation at ${\textrm{SNR}}=10$dB. Since the image regeneration tasks are the same at both SNRs, and thus the required computations are also the same, we can conclude that perceiving the meaning of the noisy received signal at the decoder demands significant computational complexity. Our findings suggest that the decoder plays a more critical role than the encoder in deepJSCC performance, especially when SNR is low. This highlights the need to reconsider the conventional symmetric deepJSCC design, where the encoder and decoder typically have symmetric structures and computational resources [5, 7, 12]. We show that designing a larger decoder than the encoder is effective in the low SNR regime, even with a simple vanilla implementation, in the ablation studies at Section IV-C.
@@
TABLE II: The ablation study for verifying the effectiveness of our proposed methods and deepJSCC complexity analysis under AWGN channel and CPP=$1/32$. PSNR ($\mathrm{dB}$) values of each SNR, and GFLOPs and Memory (MB) usages are evaluated.
SNR 1dB 4dB 7dB 10dB GFLOPs Memory (MB)
SwinJSCC [7] 24.01 24.89 25.48 25.90 20.01 71.21
LAJSCC 23.85 24.91 25.84 26.56 11.33 5.21
FAJSCC 23.96 25.15 26.07 26.86 17.53 13.47
FAJSCC w/ PGB 23.99 25.16 26.09 26.90 17.53 13.47
FAJSCC w/o AT 23.87 25.04 26.18 27.02 18.87 14.49
FAJSCC w/o FD 24.09 25.10 25.88 26.71 16.84 13.32
FAJSCC w/o LA 23.98 24.97 25.80 26.55 17.47 13.11
LAFAJSCC 24.05 25.09 25.80 26.53 15.00 9.34
FALAJSCC 23.91 24.94 25.82 26.65 15.10 9.34

### v4

@@ III-D Training and Loss Function Design
$z^{l}_{{\textrm{soft}}}(i,j)$ $=\frac{\exp{\left(\log\pi^{l}(i,j)+g^{l}(i,j)\right)/\tau}}{\sum_{m=1}^{2}\exp{\left(\log\pi^{m}(i,j)+g^{m}(i,j)\right)/\tau}},$ (12)
$\mathbf{z}_{{\textrm{hard}}}(i,j)$ $={\textrm{one\_hot}}\left(\operatorname*{arg\,max}_{l}\left[\log\pi^{l}(i,j)+g^{l}(i,j)\right]\right).$ (13)
@@
$\mathbf{z}_{{\textrm{ST}}}=\mathbf{z}^{1}_{{\textrm{hard}}}+\mathbf{z}^{1}_{{\textrm{soft}}}-{\textrm{stop\_grad}}(\mathbf{z}^{1}_{{\textrm{soft}}}).$ (14)
@@
$\mathbf{z}_{{\textrm{out}}}={\textrm{self\_attention}}({\textrm{deformation}}(\mathbf{z}_{{\textrm{ST}}}\odot_{\textrm{w}}\mathbf{z}_{{\textrm{in}}})).$ (15)
@@
$\bar{\mathbf{z}}^{e}_{{\textrm{ST}}}$ $=\frac{1}{N}\sum_{i=1}^{N}\bar{\mathbf{z}}^{e,i}_{{\textrm{ST}}},~~\bar{\mathbf{z}}^{d}_{{\textrm{ST}}}=\frac{1}{N}\sum_{i=1}^{N}\bar{\mathbf{z}}^{d,i}_{{\textrm{ST}}},~~0<\gamma_{{\textrm{train}}}<1,$ (17)
@@
Incorporating Gumbel–Softmax sampling and the regularization term in Eq. 16 (3.62 ms per image) introduces only a marginal training overhead, compared to training with $\mathcal{L}_{{\textrm{distortion}}}$ alone using random sampling without the window-importance head of Figure 3 (3.43 ms per image). In addition, the convergence behavior obtained with Eq. 16 exhibits a similar trend to that of previous deepJSCC methods [5, 6, 7, 29]. Furthermore, setting $\gamma_{{\textrm{train}}}\in[0.4,0.6]$ does not lead to noticeable performance differences. However, excessively small or large values of $\gamma_{{\textrm{train}}}$ should be avoided. A too-small $\gamma_{{\textrm{train}}}$ leads to insufficient training of the selective deformable self-attention block, whereas a too-large $\gamma_{{\textrm{train}}}$ results in inadequate training of the window-importance head, i.e., all features are always selected as important. Note that $\gamma_{{\textrm{train}}}$ is a training hyperparameter, similar in role to the learning rate.
@@ IV-A Experimental Setting
TABLE I: Architecture Summary Table for FAJSCC.
The number of stages ($L$) 44
The number of FA blocks per stage ($n_{i}$) $[2,2,2,2]$
Feature channel sizes ($C_{i}$) $[40,60,80,260]$
Window sizes $[8,8,8,8]$
The number of attention heads $[1,1,1,1]$
@@
To differentiate between these baselines, we refer to the initial deepJSCC [5] based on the convolution blocks and the deepJSCC based on residual convolution blocks [6] as ConvJSCC and ResJSCC, respectively. In contrast to SwinJSCC [7] that applies self-attention to all features, LICRFJSCC [29] applies self-attention only to a fixed subset of feature channels. The scaling settings for ConvJSCC and ResJSCC follow those in [5], while the scaling configuration for SwinJSCC is determined via an exhaustive search to achieve the best possible performance under comparable GFLOPs constraints. LICRFJSCC, LAJSCC, and FAJSCC scale settings are based on SwinJSCC. The architecture summary of FAJSCC is provided in Table I. In addition, the detailed implementation code is available at github.com/hansung-choi/FAJSCCv2.
@@
Datasets and Training To train models, we use the DIV2K training dataset [54], which consists of $800$ high-resolution 2K images. To maintain consistent image resolution within each batch and prevent out-of-memory errors, we randomly crop images to $128\times 128$ during training, with a batch size of $32$. The image distortion loss $\mathcal{L}_{{\textrm{distortion}}}$ is set to MSE and $1-{\textrm{SSIM}}$ when evaluating models for PSNR and SSIM, respectively. Our FAJSCC is trained to minimize $\mathcal{L}_{{\textrm{FA}}}$ in Eq. (16) while others are trained to minimize $\mathcal{L}_{{\textrm{distortion}}}$. The hyperparameter settings for $\mathcal{L}_{{\textrm{FA}}}$ are $\eta=1.0$, $\gamma_{{\textrm{train}}}=0.5$ , and $\tau=1$. Note that $\gamma_{{\textrm{train}}}$ is the target average importance ratio in the training phase; the importance ratio during the test phase could be different, e.g., Figure 12. All models are trained using Adam optimizer with a learning rate $0.0001$ for $200$ epochs. Experiments were conducted on a system equipped with dual Intel Xeon E5-2650 v3 CPUs (10 cores in total) and NVIDIA Tesla P100 GPU with 16 GB memory.
We consider four CPP values $\frac{1}{12},\frac{1}{16},\frac{1}{24},\frac{1}{32}]$ and four SNR values [1,4,7,10][1,4,7,10]dB under AWGN and fast Rayleigh fading channels. Except for SNR robustness experiments, the training SNR is the same as the test SNR. We evaluated deepJSCC models trained with $5$ different random seeds. For tables, values are reported as ${\textrm{mean}}\pm{\textrm{std}}$, while entries with $0.00$ std (e.g., model storage size) are omitted. For efficiency comparison plots, std values are described with vertical lines. For other plots, we omitted the standard deviation to maintain visual clarity and avoid excessive and repeated clutter.
Test results are obtained from the DIV2K validation dataset [54] and the Kodak dataset [55]. The heights and widths of the DIV2K validation images range from $1068$ to $2040$ pixels. To preserve the original spatial characteristics as closely as possible, while aligning with the SwinJSCC’s input resolution constraint, each image is center-cropped to the nearest multiple of $128$. The Kodak dataset consists of images with resolutions of $512\times 768$ or $768\times 512$. Owing to the variability in image resolutions, these datasets enable a realistic evaluation of FAJSCC and LAJSCC in practical applications, such as monitoring scenarios.
@@
Figure 6: Efficiency comparison for various model sizes with respect to computational burden (GFLOPs) and model storage size (MB) for DIV22K dataset.
TABLE II: Comparison for deployment costs of deepJSCCs under AWGN channel, SNR=10dB10\ $\mathrm${dB}, and CPP=$1/32$ for Kodak dataset.
Metric PSNR (dB) Latency (ms/image) GFLOPs Peak Memory (MB) Model Storage Size (MB)
[…]
TABLE III: Comparison for deployment costs of deepJSCCs under AWGN channel, SNR=10dB10\ $\mathrm${dB}, and CPP=$1/32$ for DIV22K dataset.
Metric PSNR (dB) Latency (ms/image) GFLOPs Peak Memory (MB) Model Storage Size (MB)
[…]
@@ IV-B Main Result
Architecture Efficiency:
To verify the performance gain of our model, we provide an efficiency comparison for various model sizes in Figure 6. To match the computational resource usage for comparison, we implement smaller and larger versions of SwinJSCC, LAJSCC, and FAJSCC by decreasing and increasing feature channel sizes by about $50\%$. In the case of ConvJSCC and ResJSCC, we increase the feature channel size by factors of $2$ and $4$ to construct larger models.
Figure 6 shows the tradeoff between computational block complexity (GFLOPs and memory usage) and PSNR performance for various JSCC schemes. Most existing works, such as ConvJSCC [5], ResJSCC [6], and SwinJSCC [7], primarily aim to improve PSNR. These performance gains are achieved by employing heavier computational blocks, which consequently incur substantially higher computational burden (GFLOPs) and model storage size (MB). In particular, SwinJSCC demands approximately an order of magnitude more model storage size than ConvJSCC and ResJSCC, and imposes a significantly higher computational burden for the same feature scales due to its reliance on heavy multi-head self-attention mechanisms. This increased requirement gives a large burden to practical deployment, especially for IoT devices. To address this issue, recent works decrease computational burden at the expense of performance degradation. For example, LICRFJSCC [29] applies advanced self-attention only to a fixed subset of feature channels without considering feature importance. Although this method reduces the computational burden, it also compromises performance by shrinking the feature representation space.
@@
Deployment Cost for Kodak dataset: For our NVIDIA Tesla P100 GPU, the models show no computation and memory traffic bottleneck for $512\times 768$ Kodak images. When peak memory is not high compared to the GPU memory bandwidth, the GPU computation follows the next structure.
$\mathrm{Memory}\rightarrow\mathrm{Compute}\rightarrow\mathrm{Compute}\rightarrow\mathrm{Store}$
In this case, the overall latency is not mainly affected by the memory traffic. Moreover, if the required GFLOPs of each block’s forward pass in the model are also not high compared to the GPU computation ability, the overall latency is also not mainly affected by the total GFLOPs. Table II precisely shows this relation. For example, ResJSCC ($5.60$ GFLOPs) has only $28.44\%$ larger GFLOPs than ConvJSCC ($4.36$ GFLOPs), but the latency is more than twice. Moreover, although GFLOPs and peak memory usage of SwinJSCC ($20.00$ GFLOPs, $493.38$ MB) are larger than our FAJSCC ($17.52$ GFLOPs, $303.86$ MB), SwinJSCC ($99.15$ ms/image) is faster than FAJSCC ($112.01$ ms/image) by about $12.97\%$.
When neither computational capacity nor memory bandwidth is a bottleneck, the overall latency is largely determined by the total number of layers in the model, as both computationally intensive and lightweight operations can be executed concurrently within the GPU’s processing capability. Table II follows this case. Since this case does not give any burden to GPU, the deployment costs are not that important. However, as image resolution increases, deployment costs become more important, as shown in the next analysis.
Deployment Cost for DIV22K dataset: For our NVIDIA Tesla P100 GPU, previous models using self-attention show a memory traffic bottleneck for high-resolution 22K DIV22K images, e.g., $1536\times 2048$ resolution. Within each window for window-based self-attention, the computation of $QK^{\top}$, followed by softmax normalization and the subsequent multiplication with $V$, requires the formation and storage of intermediate tensors. Each window independently generates attention score matrices and intermediate activations, resulting in repeated global memory reads and writes. This memory-intensive behavior significantly increases off-chip memory traffic. In this case, the GPU computation follows the next structure.
$\mathrm{Load}\rightarrow\mathrm{wait}\rightarrow\mathrm{Compute}\rightarrow\mathrm{Store}\rightarrow\mathrm{wait}\rightarrow\mathrm{Load}$
These repeated loading and waiting lead to excessive latency. For example, in Table III, huge ConvJSCC model requires $502.05$ GFLOPs and small SwinJSCC only requires $88.33$ GFLOPs. However, huge ConvJSCC ($130.48$ ms/image) is much faster than small SwinJSCC ($690.25$ ms/image) due to the small memory traffic.
The primary advantage of FAJSCC over SwinJSCC arises from the reduction in the number of processed windows via the importance ratio $\gamma$. Only half of the windows are selected for computation in FAJSCC in Table III. Consequently, compared to SwinJSCC, memory traffic is reduced significantly, and then, FAJSCC ($355.74$ ms/image) achieves less than half the latency of SwinJSCC ($704.51$ ms/image).
Figure 7: PSNR results under different channel and CPP environments for DIV22K dataset.
Figure 8: SSIM results under different channel and CPP environments for DIV22K dataset.
Figure 9: PSNR results of models trained at fixed SNR for DIV22K dataset. The first row presents the results for models trained at an SNR of $4$ dB, while the second row corresponds to models trained at an SNR of $7$ dB.
Figure 10: PSNR results of models trained at randomly sampled SNRs from [$1,4,7,10$] for DIV22K dataset.
Figure 11: PSNR results of models under the fast Rayleigh fading channel with estimated fading coefficients for DIV22K dataset.
Figure 12: Performance comparison for various importance ratios under different channel noises for DIV22K dataset. Varying $\gamma_{e}$: Only the importance ratio for the encoder’s FA blocks is changed. $\gamma_{e}=\gamma_{d}$: The importance ratios for the encoder’s and decoder’s FA blocks are changed together with the same value. Varying $\gamma_{d}$: Only the importance ratio for the decoder’s FA blocks is changed.
@@
PSNR and SSIM Results: Figures 7 and 8 show the PSNR and SSIM performance results across various communication environments. In these results, the models are trained at a fixed SNR and tested at the same SNR. Our LAJSCC and FAJSCC outperform the previous models for different performance metrics under various communication environments. This performance gain primarily comes from our dimension-specialized computation approaches: (1) Using only axis-dimension specialized operations (LAJSCC) shows higher performance than the recent SOTA SwinJSCC in most communication environments, while requiring computational resources about half and $14$ times smaller model storage size. This is a notable achievement, as even the recent SOTA lightweight model LICRFJSCC [29] reduces computational resources by shrinking feature sizes when applying self-attention, which inherently limits its performance, making it impossible to reach that of SwinJSCC. (2) Upon the axis-dimension specialized operations, our proposed selective deformable self-attention (FAJSCC) even more enhances the image transmission performance by utilizing deformation and feature importance. These advantages lead FAJSCC to outperform SwinJSCC while using smaller computational and memory resources.
SNR Robustness: Figure 9 shows the results under a single-SNR training and multi-SNR testing scenario. The first row presents the results for models trained at an SNR of $4$ dB, while the second row corresponds to models trained at an SNR of $7$ dB. On the other hand, Figure 10 shows the results of models trained at randomly sampled SNRs from [$1,4,7,10$]. In these results, our FAJSCC shows the best performance under various communication environments, and our LAJSCC shows better or comparable performance to previous deepJSCC models. Thus, our models are more robust than others under varying SNR environments.
Rayleigh Coefficient Robustness: In the above experiments on fast Rayleigh fading channels, we assumed perfect CSI only at the receiver side, i.e., the fast Rayleigh fading coefficient $\mathbf{h}$ is known. However, in real communication, the Rayleigh fading coefficient is estimated via a pilot signal. The relation between the $i$-th true Rayleigh fading coefficient $h_{i}$ and the $i$-th estimated Rayleigh fading coefficient $\hat{h}_{i}$ is as follows [56].
$\hat{h}_{i}=h_{i}+\psi_{i},~\psi_{i}\sim\mathcal{CN}(0,\sigma_{\psi}^{2}),~i\in[1:k]$ (18)
where $\psi_{i}$ is a random estimation error. When $h_{i}\sim\mathcal{CN}(0,\sigma_{\psi}^{2})$, current estimation technologies can achieve $\sigma_{\psi}^{2}=0.03$ to $\sigma_{\psi}^{2}=0.01$ [57]. To evaluate the robustness of deepJSCCs for challenging constraints, we set $\sigma_{\psi}^{2}=0.03$, which can be achieved when the SNR of the pilot signal is 5-5 dB.
Figure 11 presents the PSNR performance under the fast Rayleigh fading channel with MMSE equalization based on the estimated Rayleigh fading coefficient $\hat{\mathbf{h}}$. As observed in the figure, FAJSCC consistently achieves the best performance across varying SNRs and CPPs, showing robust superiority over the baseline models. LAJSCC also demonstrates strong performance and outperforms previous models under most SNR and CPP configurations.
@@
Computation Resource Adjustment: Our FAJSCC is the first deepJSCC model capable of independently adjusting computational resources of the transmitter and receiver. In previous computational resource-adjustable deepJSCC models, the transceiver’s computational complexity adjustments are constrained by the counterpart communication side [30, 34, 39]. As a result, such independent computational resource adjustment has not been possible in previous models. On the contrary, FAJSCC achieves this adjustment by independently varying the importance ratio $\gamma_{e}$ for the encoder’s FA blocks and $\gamma_{d}$ for the decoder’s FA blocks, without needing to agree with each other’s value. Since available computational resources for the encoder and decoder may vary over time, this flexible adjustment mechanism enhances the practicality of FAJSCC for real-world communication applications.
@@
Complexity Demands at Encoder and Decoder: Recalling that deepJSCC involves the joint optimization of source coding and channel coding, we can further investigate the image reconstruction performance from three different perspectives. The first perspective is the encoder for source and channel coding where the extracted features should be both concise and robust to channel noise. The second perspective is the decoder for source coding where images are regenerated. The final perspective is the decoder for channel coding, which involves perceiving the meaning of the noisy received signal. By considering these components, our experiments with varying values of $\gamma_{e}$ and $\gamma_{d}$ reveal the relationship between computational complexity and the contribution of each component to reconstruction performance. In previous computational resource-adjustable deepJSCC models, the transceiver’s computational complexity adjustments are constrained by the counterpart communication side [30, 34, 39]. As a result, such a discussion has not been possible in previous models.
First, when varying $\gamma_{e}$ (triangles in Figure 12), no significant change in PSNR is observed at both ${\textrm{SNR}}=1$dB (low) and ${\textrm{SNR}}=10$dB (high). This indicates that robust and concise feature extraction in the encoder does not require substantial computational resources, and adjusting the encoder’s computational budget has minimal impact on overall communication performance. On the other hand, when varying $\gamma_{d}$ (crosses in Figure 12), PSNR exhibits different behaviors depending on the SNR. The PSNR variation at ${\textrm{SNR}}=1$dB is approximately 0.280.28dB, which is about three times of the 0.090.09dB variation at ${\textrm{SNR}}=10$dB. Since the image regeneration tasks are the same at both SNRs, and thus the required computations are also the same, we can conclude that perceiving the meaning of the noisy received signal at the decoder demands significant computational complexity.
Our findings suggest that the decoder plays a more critical role than the encoder in deepJSCC performance, especially when SNR is low. This highlights the need to reconsider the conventional symmetric deepJSCC design, where the encoder and decoder typically have symmetric structures and computational resources [5, 7, 12]. We show that designing a larger decoder than the encoder is effective in the low SNR regime, even with a simple vanilla implementation, in the ablation studies at Section IV-C.
Figure 13: The first row shows the transmitted image at ${\textrm{CPP}}=\frac{1}{12}$, ${\textrm{SNR}}=10$dB under the AWGN channel, and the third row shows the transmitted image at ${\textrm{CPP}}=\frac{1}{12}$, ${\textrm{SNR}}=1$dB under the AWGN channel for Kodak dataset. The second and fourth rows show the detailed results of the green box in the original images of the first and third rows.
@@
TABLE IV: The ablation study for verifying the effectiveness of our proposed methods and deepJSCC complexity analysis under AWGN channel and CPP=$1/32$. PSNR ($\mathrm{dB}$) values of each SNR, and GFLOPs and Memory (MB) usages are evaluated for DIV22K dataset.
PSNR at SNR=11dB PSNR at SNR=44dB PSNR at SNR=77dB PSNR at SNR=1010dB Latency (ms/image) GFLOPs Peak Memory (MB) Model Storage Size (MB)
[…]